OpenStudy (jskinny0831):

The half-life of a substance is how long it takes for half of the substance to decay or become harmless (for certain radioactive materials). The half-life of a substance is 8.6 days and there is an amount equal to 15 grams now. What is the expression for the amount A(t) that remains after t days, and what is the amount of the substance remaining (rounded to the nearest tenth) after 37 days? Hint: The exponential equation for half-life is A(t) = A0(0.5)t/H, where A(t) is the final amount remaining, A0 is the initial amount, t is time, and H is the half-life. A(t) = 15(0.5)8.6t, 0.0 gram rema

OpenStudy (jskinny0831):

A. A(t) = 15(0.5)8.6t, 0.0 gram remaining B. A(t) = 15(0.5)t/8.6, 0.8 gram remaining C. A(t) = 8.6(15)(0.5)t, 0.0 gram remaining D. A(t) = 15(0.5)8.6/t, 12.8 grams remaining

OpenStudy (jskinny0831):

someone please help

OpenStudy (jskinny0831):

@ageta do you know it?

OpenStudy (ageta):

i learned it long time ago let me try to remember it

OpenStudy (jskinny0831):

okay thank you

OpenStudy (jskinny0831):

did you get it yet? @ageta

OpenStudy (wolf1728):

Here's my formulas - they are written in English as opposed to exponential equations.